OpenStudy (anonymous):

3(x - 7) - x = 2x - 21

OpenStudy (anonymous):

3(x) - 3(7)= 2x-21

OpenStudy (anonymous):

3x-21=2x-21

OpenStudy (anonymous):

3x-21-2x=2x-2x-21

OpenStudy (anonymous):

x-21=-21

OpenStudy (anonymous):

what is the final answer? x= ?

OpenStudy (anonymous):

@glubglubglub are you there?